OpenStudy (anonymous):

a number cube is tossed 24 times and lands on 1 three times and on 2 four times. find the experimental probability of landing on 1.

OpenStudy (vjt):

Mutiple choice?

OpenStudy (anonymous):

no its not multiple choice

Directrix (directrix):

Of the 24 tosses, the die landed on "1" three times. The experiemental probability = 3/24 = 1/8